# Alexandra Ciobotaru

### 11 Mar 2026 v3

@@ Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language. However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors. A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys. This method was employed successfully for English, a language with a lot of psychological NLP resources. However, this cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have created the first open-source corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community. The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
@@
Keywords mental health $\cdot$ Romanian $\cdot$ emotion detection $\cdot$ LIWC $\cdot$ topic modeling
1 Introduction
@@
Mental health research has increasingly used data from social media platforms, which often present a curated view of individuals’ personal lives. While social media can provide valuable insights for mental health studies [24, 30], there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences [32]. Another important source for mental health research is clinical data, which typically contains sensitive information and is not publicly accessible.
@@
Most research on mental health is conducted using data in English, which may overlook important nuances present in other languages. Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures [8, 30]. For example, the use of the first-person pronoun "I" has been identified as a strong predictor of depression [31], and this feature is frequently analyzed in English-based research. However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages [30]. While the pronoun "I" is often used as an indicator of depression in English, its usage requires special consideration in other languages due to linguistic differences. For instance, previous research has shown that this association has not been observed among Romanian speakers [36]. Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun ”I” in everyday communication.
@@
Despite the availability of some data in the Romanian language related to emotions, such as the RED dataset [10], REDv2 [9], and the BRIGHTER multilingual dataset [27], which includes Romanian data, there is a scarcity of mental health-related resources. Very few clinical data sources [36] and posts from medical forums [7, 12] exist, and those that do are not publicly accessible.
@@
This study aims to address this gap by creating the first publicly available dataset in Romanian, specifically designed to analyze emotional expressions in the context of depression and anxiety. By analyzing the linguistic characteristics and emotional patterns in Romanian, our research will contribute to a deeper understanding of mental health expressions in languages other than English. Through this initiative, we seek to promote more inclusive research in mental health, paving the way for targeted interventions and support across diverse linguistic communities.
2 Related Work
@@
Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language. [39] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency. [37] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses. [11] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies. This study was also done in English. [33] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods. They analyzed their 1106 responses and found various differences in semantics on the two study groups.
@@
In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: [29], [40], [2]. [4] combined LIWC features [6], sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words. [21] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales. Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
@@
3 Data Collection
As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection [5]. The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions, designed to gather extensive written responses. Out of these six questions, three touch on positive topics and three on negative topics. The survey was written in Romanian for Romanian speakers only. We further present the six open-ended questions, in both English and Romanian:
@@ 3 Data Collection
Usually, psychological corpora for NLP are gathered from social media through searching for different keywords in online posts that could indicate depression and anxiety. However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate. Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 [23], and another assessing their risk of anxiety using the GAD-7 [35]. Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3. The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
@@
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians - presidential elections and the tax increase burden generated by the country’s new leadership.
4 Methodology
Our research consists of gathering insights from PsihoRo descriptive statistics in correlation to depression and anxiety respondents’ tendencies, for three main population groups which were computed based on the severity cut-offs. According to the literature in psychology [22], a cut-off score of 10 or higher on the PHQ-9 questionnaire is used to identify the presence of potential major depressive disorder, prompting further assessment. For the GAD-7 questionnaire, also a cut-off score of 10 has been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation. [34].
Based on these rules, three population groups were derived from our dataset:
@@ 4 Methodology
Linguistic Inquiry and Word Count (LIWC) [6] is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research [31, 24, 17, 36]. The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the software are also available in several languages.
In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) [12]. The 2015 version of LIWC includes 98 categories that relate to linguistic dimensions and grammar, as well as psychological processes, including affective, cognitive, social, perceptual, and biological processes.
@@
Previous research has indicated that individuals with mental health issues tend to exhibit specific language patterns. These patterns include a higher frequency of using the personal pronoun "I" [31], increased use of words with negative connotations [14], and a tendency to use words with absolute meanings [1]. These findings highlight the importance of language analysis in understanding mental health and motivate us to use LIWC to examine the language patterns of the participants. After checking for differences in LIWC categories in relation to the PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power in classification. To this end, we employed a LightGBM classifier [20] and trained it on the results with LIWC to predict a binary label for depression and anxiety, based on the cutoff score of 10. The model achieved an accuracy of 0.85 in classifying depression and 0.83 in classifying anxiety. To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26].
@@
To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT [13] on the REDv2 dataset [9], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers. We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
@@
Another interesting analysis was topic modeling. To identify the main themes in the PsihoRo corpus we used the stm R package. First, the texts were tokenized, lowercased, and cleaned by removing punctuation, numbers, symbols and stopwords. Next, we created a Document-Feature Matrix (DFM) to represent term frequencies across texts. Also, to enhance model performance and reduce noise, we excluded rare terms which appeared in fewer than 3 documents and the very common terms, which appeared in more than 90% of the texts. Further, the cleaned DFM was transformed into the format necessary for the Structural Topic Model (STM) using which we trained the model for K = 5 topics, with the Spectral initialization method.
5 Experiments and Results
5.1 PsihoRo Descriptive Statistics
@@
In Table 1 we show mean, median (50% quantile), min, max, 25% and 75% cuantiles, and the standard deviation for both anxiety and depression scores. It can be seen that the maximum value for GAD-7 (anxiety) is 21, which is the maximum which can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
Table 1: Descriptive statistics for PsihoRo
@@
In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores. The data indicate that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
@@
The Pearson correlation coefficient (r) between PHQ-9 and GAD-7 total score is 0.761 (p<0.001), which means that anxiety and depression are strongly correlated in our data. This result shows as well in Figure 2 and is in concordance with the results of several other works: [19], [16], [18].
@@
5.2 LIWC
We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7. In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores. Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite. Asterisks indicate significance (*p < .05, **p < .01).
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination. However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety [28]. Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression ([15]). Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
Table 2: Correlations between PHQ-9, GAD-7, and LIWC dimensions
LIWC Dimension Examples PHQ-9 GAD-7
I I, me, mine 0.018 -0.046
We we, us, our 0.013 -0.039
You you, your 0.006 0.084
Verb eat, come, carry -0.087 -0.091
Past focus ago, did, talked -0.064 -0.075
Present focus today, is, now -0.019 -0.034
Future focus may, will, soon -0.027 -0.055
Adjectives free, happy, long 0.118 0.12
Negations no, not, never -.156* -.167*
Positive emotion love, nice, sweet -0.122 -0.057
Negative emotion hurt, ugly, nasty 0.028 0.028
Anxiety worried, fearful -0.042 0.01
Anger hate, kill, annoyed 0.107 0.105
Sadness crying, grief, sad 0.074 0.045
Family daughter, dad, aunt 0.079 .145*
Friends buddy, neighbor -0.026 0.043
Female references girl, her, mom 0.061 0.074
Male references boy, his, dad .156* .177*
Comparatives greater, best, after .167* 0.081
Causation because, effect -.186** -.180**
Tentative maybe, perhaps .204** .213**
Insight think, know -0.075 -0.101
Assent agree, OK, yes 0.009 0.065
Perceptual processes look, heard, feeling 0.009 0.033
Body cheek, hands, spit .200** .180**
Health clinic, flu, pill -0.051 0.035
Achievement win, success, better -.163* -.192**
Leisure cook, chat, movie -.145* -.169*
Death bury, coffin, kill .138* 0.029
@@ 5.2 LIWC
To better understand the differences between higher and lower symptom scores for depression and anxiety, we conducted an additional analysis of the LIWC categories. According to the cutoff values for both the PHQ-9 and GAD-7 assessment tools, we define a low PHQ score as anything less than 10, while a score greater than 10 is classified as a high PHQ-9 score. In Figure 3, the goal was to identify linguistic markers that differentiate individuals with higher vs. lower symptom scores. Statistically significant differences are indicated with an asterisk (*), as determined by the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores, while negative values indicate greater use among low-symptom participants. The analysis produced results consistent with the correlation analysis, particularly regarding Achievement and tentativeness in relation to depression and anxiety. However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology [36]. In addition, the language of participants with high GAD-7 scores showed a higher frequency of words in the Interrogatives category, indicating a tendency to question or seek clarity.

### 12 Mar 2026 v4

# PsihoRo: Depression and Anxiety Romanian Text Corpus

@@ Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, identify patterns related to mental health problems and analyze emotional language. However, collecting accurate mental health data from social media can be challenging due to the assumptions made by data collectors. A more effective approach involves gathering data through open-ended questions and then assessing participants’ mental health status using self-report screening surveys. This method was successfully employed for English, a language with a lot of psychological NLP resources. However, the same cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have collected the first open-source corpus focused on depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Although the PsihoRo corpus contains texts from only 205 respondents, it represents an important first step toward understanding and analyzing mental health issues within the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection, and topic modeling to identify the most important features of this newly introduced resource for the NLP community. The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
@@
Keywords: mental health, Romanian, emotion detection, LIWC, topic modeling
[…]
PsihoRo: Depression and Anxiety Romanian Text Corpus
Alexandra Ciobotaru1,2, Ana-Maria Bucur3, Liviu P. Dinu1,4
1University of Bucharest, 2Druid AI, 3Università della Svizzera italiana,
4Human Language Technologies Research Center
1,490 Panduri Street, Bucharest, Romania
21-5 Costache Negri Street, Bucharest, Romania,
3Via Buffi 13, Lugano, Switzerland
alexandra.zbaganu@s.unibuc.ro, ana-maria.bucur-cosma@usi.ch, ldinu@fmi.unibuc.ro
Abstract content
1.  Introduction
@@
Mental health research has increasingly used data from social media platforms, which often present a curated view of individuals’ personal lives. While social media can provide valuable insights for mental health studies Liu et al. (2022); Rai et al. (2024); Bucur et al. (2025), there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences Salopek and Eastin (2024). Another important source for mental health research is clinical data, which typically contains sensitive information and is often not publicly accessible Puspo et al. (2026).
@@
Most research on mental health is conducted using data in English, which may overlook important nuances present in other languages. Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures Rai et al. (2024); Bucur et al. (2026). For example, the use of the first-person pronoun “I” has been identified as a strong predictor of depression Rude et al. (2004), and this feature is frequently analyzed in English-based research. However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages Rai et al. (2024). While the pronoun “I” is often used as an indicator of depression in English, its use in other languages requires special consideration due to linguistic differences. For instance, previous research has shown that this association has not been observed among Romanian speakers Trifu et al. (2024). Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun “I” in everyday communication.
@@
Despite the availability of data in the Romanian language related to emotions, such as the RED dataset Ciobotaru and Dinu (2021), REDv2 Ciobotaru et al. (2022), and the BRIGHTER multilingual dataset Muhammad et al. (2025), which includes Romanian data, there is a scarcity of mental health-related resources. Very few clinical data sources Trifu et al. (2024) and posts from medical forums Briciu and Lupea (2018); Dudău and Sava (2022) exist, but unfortunately are not publicly accessible.
@@
This study aims to address this gap by creating the first publicly available dataset in Romanian, specifically designed to analyze emotional expressions in the context of depression and anxiety. By analyzing the linguistic characteristics and emotional patterns in Romanian, our research aims to contribute to a deeper understanding of mental health expressions in languages other than English. Through this initiative, we seek to promote more inclusive research in mental health, paving the way for targeted interventions and support across diverse linguistic communities.
2.  Related Work
@@
Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language. Yoon et al. (2024) created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. The authors gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency. Van der Vegt and Kleinberg (2023) created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses. Dai et al. (2022) used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual in English. Sikström et al. (2023) used open-ended questionnaires to study young and middle-aged adults’ responses regarding their own mental health, by employing NLP computational methods. The authors analyzed 1,106 responses and found semantic differences between the two study groups.
@@
In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: Rahman et al. (2024), Zhang et al. (2023), Alshouha et al. (2024). Beech et al. (2025) combined LIWC features Boyd et al. (2022), sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words. Kjell et al. (2021) collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales. Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
@@
3.  Data Collection
As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection Benton et al. (2017). The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions, designed to gather extensive written responses. Out of these six questions, three touch on positive topics and three on negative topics. The survey was written in Romanian for Romanian speakers only. We further present the six open-ended questions, in both English and Romanian:
@@ 3.  Data Collection
Usually, psychological corpora for NLP are gathered from social media through searching for different keywords in online posts that could indicate depression and anxiety. However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate. Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 Kroenke et al. (2001), and another assessing their risk of anxiety using the GAD-7 Spitzer et al. (2006). Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and each item is answered on a 4-point Likert scale: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3. The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
@@
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians–presidential elections and the tax increase burden generated by the country’s new leadership.
4.  Methodology
Our research consists of gathering insights from PsihoRo using descriptive statistics to analyze the correlation between depression and anxiety levels among respondents across three main population groups, computed based on severity cut-offs. According to the literature in psychology Kroenke et al. (2001), a PHQ-9 score of 10 or higher indicates the potential presence of major depressive disorder and prompts further assessment. For the GAD-7 questionnaire, a cut-off score of 10 has also been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation Spitzer et al. (2006).
Based on these cut-off scores, three population groups were derived from our dataset:
@@ 4.  Methodology
Linguistic Inquiry and Word Count (LIWC) Boyd et al. (2022) is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research Rude et al. (2004); Liu et al. (2022); Hur et al. (2024); Trifu et al. (2024). The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the lexicon are also available in several languages. In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) Dudău and Sava (2022). The 2015 version of LIWC includes 98 categories that relate to linguistic dimensions and grammar, as well as psychological processes, including affective, cognitive, social, perceptual, and biological processes.
@@
Previous research on the English language has indicated that individuals with mental health issues tend to exhibit specific language patterns. These patterns include a higher frequency of using the personal pronoun “I” Rude et al. (2004), increased use of words with negative connotations Fekete (2002), and a tendency to use words with absolute meanings Al-Mosaiwi and Johnstone (2018). These findings highlight the importance of language analysis for understanding mental health and motivate us to use LIWC to examine participants’ language patterns. After checking for differences in LIWC categories across PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power for classification. To this end, we employed a LightGBM classifier Ke et al. (2017) and trained it on LIWC results to predict a binary label for depression and anxiety, using a cut-off score of 10. The model achieved accuracies of 0.85 for depression and 0.83 for anxiety. To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017).
@@
To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT Dumitrescu et al. (2020) on the REDv2 dataset Ciobotaru et al. (2022), for multi-label text classification, using the Huggingface transformers pipeline. [^1] https://huggingface.co/docs/transformers We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model achieved an F1 score of 66.85% on the REDv2 test set.
@@
We also performed topic modeling to identify the main themes in the PsihoRo corpus, using the stm R package. First, the texts were tokenized, lowercased, and cleaned by removing punctuation, numbers, symbols and stopwords. Next, we created a Document-Feature Matrix (DFM) to represent term frequencies across texts. To enhance model performance and reduce noise, we excluded rare terms that appeared in fewer than 3 documents and the very common terms that appeared in more than 90% of the texts. Further, the cleaned DFM was transformed into the format required by the Structural Topic Model (STM), and we trained the model with $K$ = 5 topics using the Spectral initialization method.
5.  Experiments and Results
5.1.  PsihoRo Descriptive Statistics
@@
In Table 1, we show mean, median (50% percentile), min, max, 25% and 75% percentiles, and the standard deviation for both anxiety and depression scores. In the collected data, the maximum value for GAD-7 (anxiety) is 21, which is the maximum that can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
@@
In Figure 1, we present the distribution plots of both PHQ-9 and GAD-7 scores. The data indicate that the distributions are skewed to the left, with most observations clustered at lower values and a longer tail extending toward the higher end of the scale. This suggests that the majority of respondents experience minimal to mild risk of depression and anxiety.
@@
The Pearson correlation coefficient ($r$) between the PHQ-9 and GAD-7 total scores is 0.761 (pp<0.001), indicating that anxiety and depression are strongly correlated in our data. This finding is illustrated in Figure 2 and aligns with the results of several other studies Kalin (2020); Hoying et al. (2020); Jacobson and Newman (2014).
@@
LIWC Dimension Examples PHQ-9 GAD-7 I I, me, mine .018 -.046 We we, us, our .013 -.039 You you, your .006 .084 Verb eat, come, carry -.087 -.091 Past focus ago, did, talked -.064 -.075 Present focus today, is, now -.019 -.034 Future focus may, will, soon -.027 -.055 Adjectives free, happy, long .118 .120 Negations no, not, never -.156* -.167* Positive emotion love, nice, sweet -.122 -.057 Negative emotion hurt, ugly, nasty .028 .028 Anxiety worried, fearful -.042 .010 Anger hate, kill, annoyed .107 .105 Sadness crying, grief, sad .074 .045 Family daughter, dad, aunt .079 .145* Friends buddy, neighbor -.026 .043 Female references girl, her, mom .061 .074 Male references boy, his, dad .156* .177* Comparatives greater, best, after .167* .081 Causation because, effect -.186** -.180** Tentative maybe, perhaps .204** .213** Insight think, know -.075 -.101 Assent agree, OK, yes .009 .065 Perceptual processes look, heard, feeling .009 .033 Body cheek, hands, spit .200** .180** Health clinic, flu, pill -.051 .035 Achievement win, success, better -.163* -.192** Leisure cook, chat, movie -.145* -.169* Death bury, coffin, kill .138* .029
@@
5.2.  LIWC
We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7. In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores. Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite.